OpenStudy (loser66):

Suppose that for each in P the function y is defined by \[y (x)= \int_0^2 (x (t))^2dt\] is y a linear functional? Please, help

OpenStudy (loser66):

By definition, y is called a linear functional if it satisfy the condition \[y(\alpha x_1+\beta x_2)= \alpha y(x_1)+\beta y(x_2)\]

OpenStudy (loser66):

My work: \[y (\alpha x_1) = \int_0^2 \alpha^2(x_1(t))^2 dt~~ (1)\] I want to confirm this equation first. Am I right or \[y (\alpha x_1) = \int_0^2 \alpha(x_1(t))^2 dt ~~(2)\] which one is the right one (1) or (2) then, I am OK

OpenStudy (zarkon):

\[y (\alpha x_1) = \int_0^2 (\alpha x_1(t))^2 dt=\int_0^2 \alpha^2(x_1(t))^2 dt\]

OpenStudy (zarkon):

what is P?

OpenStudy (loser66):

Thank you @Zarkon. I can step up from here. I don't want to go so far if I have a mistake there, :) P is a polynomial

OpenStudy (loser66):

*set

OpenStudy (zarkon):

ok

OpenStudy (ybarrap):

looks good

OpenStudy (loser66):

How about \[y(x) = \int_0^1 t^2x(t) dt\]

OpenStudy (loser66):

\[y(\alpha x_1(t) = \alpha^2t^2x_1(t) dt\] right?

OpenStudy (loser66):

oh, I forgot \(\int_0^1\) in the front

OpenStudy (ybarrap):

yes. You are checking if this particular parametrized equations are linear, so yes. That's how you would check if functions scales.

OpenStudy (ybarrap):

not y(x(at)) but y(ax(t), like you did, yes.

OpenStudy (loser66):

what is difference between the 2?

OpenStudy (loser66):

oh, I got it, friend. like f (3x) \(\neq\) 3 f(x). right?

OpenStudy (ybarrap):

For example, if you ook at a line parametrized by arc length. If S(t) is arclength (parametrized) then aS(t) is \(a\) times bigger than S(t). But S(at) is not \(a\) times bigger than S(t)

OpenStudy (ybarrap):

yep, that's it

OpenStudy (loser66):

got you :)

OpenStudy (ybarrap):

kool!

OpenStudy (loser66):

hey, I still have one more, hehehe. need check y(x) = \(\dfrac{d^2x}{dt^2}|t=1\) it looks not nice at all.

OpenStudy (zarkon):

The derivative is a linear operator